OpenStudy (blackasinc):

Prove: cot x cos x + sin x = csc x I believe I start with left side: cos/ sin * cos + sin ...but what next

satellite73 (satellite73):

add

satellite73 (satellite73):

you have \[\frac{\cos^2(x)}{\sin(x)}+\sin(x)\] on the left add them

OpenStudy (blackasinc):

do i need to multiply both sides by 1/sin so they have same denominator?

OpenStudy (blackasinc):

before i add

OpenStudy (mathstudent55):

No. Just work on the left side of the identity. Multiply \(\sin x\) by \(\dfrac{\sin x}{\sin x} \) to have a common denominator. Then add the two fractions on the left side together. Then pay close attention to the numerator after you add the fractions.

OpenStudy (blackasinc):

thanks i got it!